Sunland Sports has determined the following unit costs for each of its products:

Pants Jerseys Jackets

Direct materials

$7.00 $10.00 $58.00

Direct labor

4.00 3.60 22.00

Manufacturing overhead

5.00 4.00 34.00

Total unit cost

$16.00 $17.60 $114.00

Sales price per unit

$21.00 $22.40 $179.00


On May 31, Sunland Sports’ Work in Process Inventory consisted of the following items:

Job Units Accumulated Cost

PA-1247—Pants

100 $1,150

JE-1397—Jerseys

200 2,800

JA 426—Jackets

50 4,350
$8,300


During June, a total of $267,900 in direct materials and $125,380 in direct labor costs were incurred. Units finished and sold during June were as follows:

Product Units Finished Units Sold

Pants

13,500 14,000

Jerseys

3,200 3,100

Jackets

2,500 2,500

Given that Sunland Sports uses direct labor dollars as its application base, what is the company’s predetermined overhead rate? (Round answer to 0 decimal places, e.g. 15%.)

??? %

  Calculate the total manufacturing cost for June. (Round answer to 0 decimal places, e.g. 5,275.)

???

.Calculate the Cost of Goods Manufactured for June. (Round answer to 0 decimal places, e.g. 5,275.)

???

Calculate the Ending Work in Process Inventory balance on June 30. (Round answer to 0 decimal places, e.g. 5,275.)

???

Calculate Cost of Goods Sold for June. (Round answer to 0 decimal places, e.g. 5,275.)

???

Calculate Gross Profit for June.

???

For each of the three products, state whether there are more, fewer, or the same number of finished units in Finished Goods Inventory on June 30 than there were on June 1.

Pants

select an option                                                                      No changeMoreFewer

Jerseys

select an option                                                                      FewerNo changeMore

Jackets

select an option                                                                      FewerMoreNo change
Answer #1

Answer:

Particulars Pants Jersey's Jackets Totals
Overhead cost 5 4 34 43
Direct Labor Cost (DLC) 4 3.60 22 29.6
Predetermined Overhead rate 125%
Direct Material a) 267,900
Direct Labor b) 125,380
Overhead (125,380 * 125%) c) 156,725
Total Manufacturing Cost (a+b+c) 550,005

Total Cost of Goods Manufactured = 13,500 * 16 + 3,200 * 17.60 + 2,500 * 114

= 216,000 + 56,320 + 285,000

  = 557,320

Closing Work In Process = Opening WIP + Total Manufacturing Cost - Total Cost of Goods Manufactured

= 8,300 + 550,005 - 557,320

= 985

Cost Of Goods Sold (COGS) = 14,000 * 16 + 3,100 * 17.60 + 2,500 * 114

= 224,000 + 54,560 + 285,000

= 563,560

Gross Profit = Total Sales - COGS

  = 810,940 - 563,560

= 247,380

Workings:

Total Sales = 14,000 * 21 + 3,100 * 22.40 + 2,500 * 179

= 294,000 + 69,440 + 447,500

= 810,940